1 We have an IS-LM Model:


 C = 200 + 0.25(Y-T) 

I = 150 + 0.25Y – 1000i 

G = 250 

T = 200 

,    

;  

 

Derive IS curve both mathematically and graphically

Determine the slope of IS curve (di/dY)

Identify the shift variables of the IS curve

Derive aggregate demand curve both graphically and mathematically

Find equilibrium level of output and the interest rate

Derive the following multipliers

 

Determine the impact of simultaneous increase in government expenditure and money supply (such that ) on equilibrium output and the interest rate.


1
Expert's answer
2021-04-12T07:08:05-0400

IS curve is:

Y = C + I + G = 200 + 0.25(Y - 200) + 150 + 0.25Y – 1000i + 250,

0.5Y = 550 - 1000i,

Y = 1100 - 2000i or i = 0.55 - 0.001Y.

The slope is -0.001.

The shift variables of the IS curve are: C, I, G, T, i.

AD = Y = 550 - 1000i.

We need to know additional data about the LM curve to find equilibrium level of output and the interest rate.

The expenditure multiplier is:

"m = 1\/(1 - c) = 1\/0.75 = 4\/3."

A simultaneous increase in government expenditure and money supply will increase AD, as a result the equilibrium output will increase and the interest rate will decrease.